# “How Big is Big Enough?” Adjusting Model Size in Continual Gaussian Processes

### v1

@@ Abstract
For many machine learning methods, creating a model requires setting a parameter that controls the model’s capacity before training, e.g. number of neurons in DNNs, or inducing points in GPs. Increasing capacity improves performance until all the information from the dataset is captured. After this point, computational cost keeps increasing, without improved performance. This leads to the question “How big is big enough?” We investigate this problem for Gaussian processes (single-layer neural networks) in continual learning. Here, data becomes available incrementally, and the final dataset size will therefore not be known before training, preventing the use of heuristics for setting the model size. We provide a method that automatically adjusts this, while maintaining near-optimal performance, and show that a single hyperparameter setting for our method performs well across datasets with a wide range of properties.
@@ 1 Introduction
Continual learning aims to train models when the data arrives in a stream of batches, without storing data after it has been processed, and while obtaining predictive performance that is as high as possible at each point in time [35]. Selecting the size of the model is challenging in this setting, since typical non-continual training procedures do this by trial-and-error (cross-validation) using repeated training runs, which is not possible under our requirement of not storing any data. Selecting model size is crucial, since if the model is too small, predictive performance will suffer. One solution could be to simply make all continual learning models so large, that they will always have enough capacity, regardless of what dataset and what amount of data they will be given. However, this “worst-case” strategy is wasteful of computational resources.
A more elegant solution would be to grow the size of the model adaptively as data arrives, according to the needs of the problem (see Figure 1 for an illustration). For example, if data were only ever gathered from the same region, there would be diminishing novelty in every new batch, leading to a possible halt in growth, with growth resuming once data arrives from new regions. In this paper, we investigate a principle for determining how to select the size of a model so that it is sufficient to obtain near-optimal performance, while otherwise wasting a minimal amount of computation. In other words, we seek to answer the question of “how big is big enough?” for setting the size of models throughout continual learning.
@@
Our approach benefits from Bayesian nonparametric perspectives, by separating the specification of the capacity of the model, which should be able to learn from an unbounded amount of data, and the specification of the approximation, which determines how much computational effort is needed to represent the solution for the current finite dataset. Our approach relies on the variational inducing variable approximation [41], which leads to an interpretation that our method automatically adjusts the width of a single-layer neural network. Our hope is that such perspectives will be useful for the development of adaptive deep neural networks, which would allow models to be trained on large amounts of data beyond the limitations of storage infrastructure, or learning from streams of data that come directly from interaction with an environment.
Figure 1: Top: Three consecutive batches of data for different continual learning regression scenarios. Bottom: Number of inducing points selected using the VIPS algorithm at each batch for each dataset, demonstrating different behaviour based on the problem’s demands.
@@ 2 Related Work
The most widely discussed problem in continual learning is that of catastrophic forgetting, where previously acquired knowledge is lost in favour of recent information [28, 17]. Many solutions have been proposed in the literature, such as encouraging weights to be close to values that were well-determined by past data [21, 38], storing subsets or statistics of past data to continue to train the neural network on in the future [22, 23], and approximate Bayesian methods that balance uncertainty estimates of parameters with the strength of the data [32, 36, 5]. Within continual learning, many different settings have been investigated, which vary in difficulty [11]. Across these tasks, the gap in performance to a full-batch training procedure therefore also varies, but despite progress, some gap in performance remains.
@@
Bayesian continual learning methods have been developed because the posterior given past data becomes the prior for future data, making the posterior a sufficient quantity to estimate [30]. For the special case of linear-in-the-parameters regression models, the posterior and updates can be calculated in closed form, leading to continual learning giving exactly the same result as full-batch training. In most cases (e.g. for neural networks), the posterior cannot be found exactly, leading to the aforementioned methods [32, 36, 5] that focus on finding an approximation to the posterior and using this as the sufficient quantity.
@@
Even with a perfect solution to catastrophic forgetting (e.g. in the case of linear-in-the-parameters regression models), continual learning methods face the additional difficulty of ensuring that models have sufficient capacity to accommodate the continuously arriving information. In continual learning, it is particularly difficult to determine a fixed size for the model, since the number of data or tasks are not yet known, and selecting a model that is too small can significantly hurt performance. To improve over a fixed model size, methods can be made to grow with the data size. For example, Rusu et al. [37] extend hidden representations by a fixed amount for each new batch of data that arrives, and allows the weights of the extended representation to depend on the representation of all previous tasks. Yoon et al. [45] argue that extension by a fixed amount is wasteful and should instead be data dependent, specifically by copying neurons if their value changes too much, and adding new neurons if the training loss doesn’t reach a particular threshold. Kessler et al. [20] propose to use the Indian Buffet Process as a more principled way to regularise how fast new weights are added with tasks. While the data dependence that both these methods introduce is necessary to prevent computational waste, both methods have hyperparameters that need to be tuned to dataset characteristics, which is difficult when the dataset characteristics are not known at the start of training.
@@
Growing model capacity with dataset size was one of the main justifications for research into (Bayesian) non-parametric models [15, 16]. This approach defines models with infinite capacity, with Bayesian inference naturally using an appropriate finite capacity to make predictions, with finite compute budgets. Gaussian processes (GPs) [34] are the most common Bayesian non-parametric model for supervised learning, and are equivalent to infinitely-wide deep neural networks [31, 27] and linear-in-the-parameters models with an infinite feature space. Their infinite capacity allows them to recover functions perfectly in the limit of infinite data [43], and their posterior can be computed in closed form. These two mathematical properties provide strong principles for providing high-quality solutions to both catastrophic forgetting and ensuring appropriate capacity, and therefore make GPs an excellent model for studying continual learning.
@@
However, developing a practical continual learning in GPs is not as straightforward as it is in finite dimensional linear models, because (for $N$ datapoints) the posterior requires 1) $O(N^{3})$ operations to compute it exactly, which becomes intractable for large datasets, and 2) storing the full training dataset, which breaks the requirements of continual learning. Sparse variational inducing variable methods have been proposed to solve these problems [41], by introducing a small number of $M$ inducing points that control the capacity of the posterior approximation. In certain settings, this approximation is near-exact even when $M\ll N$ [3]. This property has allowed continual learning methods to be developed for GPs that perform very closely to full-batch methods [2, 24, 4], provided $M$ is large enough.
@@
As in neural network models, selecting the capacity $M$ is an open problem, with several proposed solutions. Kapoor et al. [19] acknowledge the need for scaling the capacity with data size, and propose VAR-GP (Variational Autoregressive GP) which adds a fixed number of inducing points for every batch. However, this number may be too small, leading to poor performance, or too large, leading to wasted computation. Galy-Fajou and Opper [14] propose OIPS (online inducing point selection), which determines $M$ through a threshold on the correlation with other inducing points, which needs to be tuned based on dataset properties.
@@
In this work, we propose to instead select the capacity of the variational approximation by selecting an appropriate tolerance in the KL gap to the true posterior. This criterion works within the same computational constraints as existing GP continual learning methods, adapts the capacity to the dataset to minimise computational waste while retaining near-optimal performance. Our method has a single hyperparameter that we keep fixed to a single value, and that produces similar trade-offs across benchmark datasets with significantly different characteristics.
@@ 3.1 Sparse Variational Gaussian Processes
We consider the typical regression setting, with training data consisting of $N$ input/output pairs $\{\mathbf{x}_{n},y_{n}\}_{n=1}^{N},\mathbf{x}_{n}\in\mathbb{R}^{D},y_{n}\in\mathbb{R}$. We model $y_{n}$ by passing $\mathbf{x}_{n}$ through a function followed by additive Gaussian noise $y_{n}=f(\mathbf{x}_{n})+\epsilon_{n},\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2})$, and take a Gaussian process prior on $f\sim\mathcal{GP}(0,k_{\theta}(\cdot,\cdot))$ with zero mean, and a kernel $k$ with hyperparameters $\theta$. While the posterior (for prediction) and marginal likelihood (for finding $\theta$) can be computed in closed form [34], they have a computational cost of $O(N^{3})$ that is too high, and require all training data (or statistics greater in size) to be stored, both of which are prohibitive for continual learning. Variational inference can provide an approximation at a lower $O(NM^{2})$ computational and $O(NM)$ memory costs by selecting an approximation from a set of tractable posteriors
@@
$=\mathcal{N}\left(f(\cdot);\mathbf{k}_{\cdot\mathbf{u}}\mathbf{K}_{\bf uu}^{-1}\mathbf{m},k(\cdot,\cdot)-\mathbf{k}_{\cdot\mathbf{u}}\mathbf{K}_{\bf uu}^{-1}(\mathbf{K}_{\bf uu}-\mathbf{S})\mathbf{K}_{\bf uu}^{-1}\mathbf{k}_{\mathbf{u}\cdot}\right)\,,$ (2)
with $[\mathbf{K}_{\bf uu}]_{ij}=k(\mathbf{z}_{i},\mathbf{z}_{j})$, $[\mathbf{k}_{\cdot\mathbf{u}}]_{i}=[\mathbf{k}_{\mathbf{u}\cdot}^{T}]_{i}=k(\cdot,\mathbf{z}_{i}),\mathbf{Z}=\{\mathbf{z}_{m}\}_{m=1}^{M}$, and $q(\mathbf{u})=\mathcal{N}(\mathbf{u};\mathbf{m},\mathbf{S})$. The variational parameters $\mathbf{m},\mathbf{S},\mathbf{Z}$ and hyperparameters $\theta$ are selected by maximising the Evidence Lower Bound (ELBO). This simultaneously minimises KL gap $\mathrm{KL}[q(f)\,||\,p(f|\mathbf{y},\theta)]$ between the approximate and true GP posteriors [26, 25], and maximises an approximation to the marginal likelihood of the hyperparameters:
$\mathcal{L}_{\text{ELBO}}=\sum_{i=1}^{N}\mathbb{E}_{q(f(\mathbf{x}_{i}))}[\log p(y_{i}|f(\mathbf{x}_{i}),\theta)]-\mathrm{KL}\left[q(\mathbf{u})\,\|\,p(\mathbf{u}|\theta)\right]\;.$ (3)
The variational approximation has the desirable properties [44] of 1) providing a measure of discrepancy between the finite capacity approximation, and the true infinite capacity model, 2) resulting in arbitrarily accurate approximations if enough capacity is added [3], and 3) retaining the uncertainty quantification over the infinite number of basis functions. In this work, we will particularly rely on being able to measure the quality of the approximation to help determine how large $M$ should be.
@@ 3.2 Sparse Gaussian Processes are Equivalent to Single-Layer Neural Networks
For inner product kernels $k(\mathbf{x},\mathbf{Z})=\sigma(\mathbf{Z}\mathbf{x})$ like the arc-cosine kernel [6], the mean is equivalent to a single-layer neural network with $\mathbf{Z}$ as the input weights, and $\mathbf{K}_{\bf uu}^{-1}\mathbf{m}$ as the output weights. This construction also arises from other combinations of kernels and inter-domain inducing variables [9, 40], and has also showed equivalences between deep Gaussian processes and deep neural networks [10]. As a consequence, our method for determining the number of inducing variables needed in a sparse GP equivalently finds the number of neurons needed in a single-layer neural network.
@@ 3.3 Online Sparse Gaussian Processes
In this work, we use the extension of the sparse variational GP approximation to the continual learning case developed by Bui et al. [2]. We modified the typical derivation to 1) clarify how the online ELBO provides an estimate to the full-batch ELBO, and 2) clarify when this approximation is accurate.
In this online setting, we aim to update our posterior and hyperparameter approximations after each batch of new data $\{\mathbf{X}_{n},\mathbf{y}_{n}\}$. While we do not have access to data from older batches $\{\mathbf{X}_{o},\mathbf{y}_{o}\}$, the parameters specifying the approximate posterior $q_{o}(f)=p(f_{\neq\mathbf{a}}|\mathbf{a},\theta_{o})q_{o}(\mathbf{a})$ are passed on. This approximate posterior is constructed as in eq. (1) but with $\mathbf{a}=f(\mathbf{Z}_{o})$ and the old hyperparameters $\theta_{o}$. Given this information, online sparse GP approaches aim to obtain an approximation $q_{n}(f)$ of the posterior distribution for all observed data $p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})$, which we rewrite as:
$p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})=\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)p(\mathbf{y}_{o}|f)}{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}=\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)}{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}\frac{p(f|\mathbf{y}_{o},\theta_{o})p(\mathbf{y}_{o}|\theta_{o})}{p(f|\theta_{o})}\;.$
We denote the new variational distribution as $q_{n}(f)=p(f_{\neq\mathbf{b}}|\mathbf{b},\theta_{n})q_{n}(\mathbf{b})$ where $\mathbf{b}=f(\mathbf{Z}_{n})$ and $\theta_{n}$ is the new hyperparameter which can differ from $\theta_{o}$. The KL divergence between the exact and approximate posterior at the current batch is given by:
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]=\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}-\int q_{n}(f)\log\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)p(f|\mathbf{y}_{o},\theta_{o})}{q_{n}(f)p(f|\theta_{o})}\text{d}f\;.$
@@
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]=\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}-\int q_{n}(f)\log\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)q_{o}(f)}{q_{n}(f)p(f|\theta_{o})}\text{d}f+\Phi$ (4)
where $\Phi=-\int q_{n}(f)\log\frac{p(f|\mathbf{y}_{o},\theta_{o})}{q_{o}(f)}\text{d}f$. We cannot compute $\Phi$ due to its dependence on the exact posterior, so we drop it and use the following “online ELBO” as our training objective:
^$\widehat{\mathcal{L}}:=\int q_{n}(f)\left[\log\frac{p(\mathbf{b}|\theta_{n})q_{o}(\mathbf{a})p(\mathbf{y}_{n}|f)}{q_{n}(\mathbf{b})p(\mathbf{a}|\theta_{o})}\right]\text{d}f.$ (5)
Maximising $\widehat{\mathcal{L}}$ will accurately minimise the KL to the true posterior when $\Phi$ is small, which is the case when the old approximation is accurate, i.e. $q_{o}(f)\approx p(f|\mathbf{y}_{o},\theta_{o})$ for all values of $f$ (with $\Phi=0$ in the case of equality). In our continual learning procedure, we will keep our sequence of approximations accurate by ensuring they all have enough inducing points.
To get our final bound, we perform a change of variables for the variational distribution $q_{o}(\mathbf{a})=\mathcal{N}(\mathbf{a};\mathbf{m}_{\mathbf{a}},\mathbf{S}_{\mathbf{a}})$ to use the likelihood parametrisation [33]:
$q_{o}(\mathbf{a})=\frac{\mathcal{N}(\mathbf{a};\tilde{\mathbf{m}}_{\mathbf{a}},\mathbf{D}_{\mathbf{a}})\mathcal{N}(\mathbf{a};0,\mathbf{K}^{\prime}_{\mathbf{aa}})}{\int\mathcal{N}(\mathbf{a};\tilde{\mathbf{m}}_{\mathbf{a}},\mathbf{D}_{\mathbf{a}})\mathcal{N}(\mathbf{a};0,\mathbf{K}^{\prime}_{\mathbf{aa}})\text{d}\mathbf{a}}=\frac{l(\mathbf{a})p(\mathbf{a}\mid\theta_{o})}{\mathcal{N}(\mathbf{a};0,\mathbf{D}_{\mathbf{a}}+\mathbf{K}^{\prime}_{\mathbf{aa}})},$ (6)

### v4

# Adjusting Model Size in Continual Gaussian Processes: How Big is Big Enough?

@@ Abstract
Many machine learning models require setting a parameter that controls their size before training, e.g. number of neurons in DNNs, or inducing points in GPs. Increasing capacity typically improves performance until all the information from the dataset is captured. After this point, computational cost keeps increasing, without improved performance. This leads to the question “How big is big enough?” We investigate this problem for Gaussian processes (single-layer neural networks) in continual learning. Here, data becomes available incrementally, and the final dataset size will therefore not be known before training, preventing the use of heuristics for setting a fixed model size. We develop a method to automatically adjust model size while maintaining near-optimal performance. Our experimental procedure follows the constraint that any hyperparameters must be set without seeing dataset properties, and we show that our method performs well across diverse datasets without the need to adjust its hyperparameter, showing it requires less tuning than others.
Machine Learning, Continual Learning, Gaussian Processes, Adaptive Model Capacity, Variational Inference, Inducing Points, Sparse Approximations, Bayesian Nonparametrics
@@ 1 Introduction
Figure 1: Three continual learning scenarios with different capacity requirements. Top: Three consecutive batches for 1) a growing input space, 2) i.i.d. samples from a uniform distribution, and 3) narrow-range samples with occasional outliers. Bottom: Number of inducing points selected using the VIPS algorithm at each batch. When selecting model size with our method, we observed: 1) a linear increase, 2) after initial training, we see a halt in growth, and 3) a low model size until it encounters outliers.
Continual learning aims to train models when the data arrives in a stream of batches, without storing data after it has been processed, and while obtaining predictive performance that is as high as possible at each point in time (Ring, 1997). Selecting the size of the model is challenging in this setting, since typical non-continual training procedures do this by trial-and-error (cross-validation) using repeated training runs, which is not possible under our requirement of not storing any data. This is a crucial parameter to set well, since if the model is too small, predictive performance will suffer. One solution could be to simply make all continual learning models so large, that they will always have enough capacity, regardless of what dataset and what amount of data they will be given. However, this “worst-case” strategy is wasteful of computational resources.
A more elegant solution would be to grow the size of the model adaptively as data arrives, according to the needs of the problem (see Figure 1 for an illustration). For example, if data were only ever gathered from the same region, there would be diminishing novelty in every new batch. In this case, one would want the model size to stabilise, with growth resuming once data arrives from new regions. In this paper, we investigate a principle for determining how to select the size of a model so that it is sufficient to obtain near-optimal performance, while otherwise wasting a minimal amount of computation. In other words, we seek to answer the question of “how big is big enough?” for setting the size of models throughout continual learning.
@@
Our approach benefits from Bayesian nonparametric perspectives, by separating the specification of the capacity of the model, which should be able to learn from an unbounded amount of data, from the specification of the approximation, which determines how much computational effort is needed to represent the solution for the current finite dataset. Our approach relies on the variational inducing variable approximation of Titsias (2009), which leads to an interpretation that our method automatically adjusts the width of a single-layer neural network. Our hope is that such perspectives will be useful for the development of adaptive deep neural networks, which would allow models to be trained on large amounts of data beyond the limitations of storage infrastructure, or learning from streams of data that come directly from interaction with an environment.
@@ 2 Related Work
The foundational problem that makes training with gradient descent insufficient for continual learning is catastrophic forgetting, where previously acquired knowledge is lost in favour of recent information (McCloskey & Cohen, 1989; Goodfellow et al., 2013). Many solutions have been proposed in the literature, such as encouraging weights to be close to values that were well-determined by past data (Kirkpatrick et al., 2017; Schwarz et al., 2018), storing subsets or statistics of past data to continue to train the neural network in the future (Li & Hoiem, 2017; Lopez-Paz & Ranzato, 2017), and approximate Bayesian methods that balance uncertainty estimates of parameters with the strength of the data (Bui et al., 2017; Nguyen et al., 2018; Rudner et al., 2022; Chang et al., 2023b). Within continual learning, many different settings have been investigated, which vary in difficulty (Farquhar & Gal, 2018). Across these tasks, the gap in performance to a full-batch training procedure therefore also varies, but despite progress, some gap in performance remains.
@@
Bayesian continual learning methods have been developed because the posterior given past data becomes the prior for future data, making the posterior a sufficient quantity to estimate (Murphy, 2023, §19.7). For the special case of linear-in-the-parameters regression models, the posterior and updates can be calculated in closed form, leading to continual learning giving exactly the same result as full-batch training. In most cases (e.g. for neural networks), the posterior cannot be found exactly, necessitating the aforementioned approximations that use an approximate posterior as the sufficient quantity (Nguyen et al., 2018; Rudner et al., 2022; Chang et al., 2023b).
@@
Even with a perfect solution to catastrophic forgetting (e.g. in the case of linear-in-the-parameters regression models), continual learning methods face the additional difficulty of ensuring that models have sufficient capacity to accommodate the continuously arriving information. In continual learning, it is particularly difficult to determine a fixed size for the model, since the number of data or tasks are not yet known, and selecting a model that is too small can significantly hurt performance. Growing models with data can fix this problem. For example, Rusu et al. (2016) extend hidden representations by a fixed amount for each new batch. Yoon et al. (2018) argue that extension by a fixed amount is wasteful and should instead be data dependent, specifically by copying neurons if their value changes too much, and adding new neurons if the training loss doesn’t reach a particular threshold. Kessler et al. (2021) propose to use the Indian Buffet Process as a more principled way to regularise how fast new weights are added with tasks. Maile et al. (2022); Mitchell et al. (2023) develop similar ideas of automatically adjusting model size to the normal batch-learning setting. We agree with the aims of these papers, but note that they rely on thresholds that need to be tuned to the dataset characteristics, e.g. tolerances on the training error which depend on the amount of intrinsic noise in the dataset. We seek a unifying training principle that is broadly applicable to a wide range of datasets, without knowing their properties a priori.
@@
Growing model capacity with dataset size was one of the main justifications for research into (Bayesian) non-parametric models (Ghahramani, 2013, 2015). This approach defines models with infinite capacity, with Bayesian inference naturally using an appropriate finite capacity to make predictions, with finite compute budgets. Gaussian processes (GPs) (Rasmussen & Williams, 2005) are the most common Bayesian non-parametric model for supervised learning, and are equivalent to infinitely-wide deep neural networks (Neal, 1996; Matthews et al., 2018) and linear-in-the-parameters models with an infinite feature space (Mercer, 1909). Their infinite capacity allows them to recover functions perfectly in the limit of infinite data (van der Vaart & van Zanten, 2008), and their posterior can be computed in closed form. These are strong principles for providing high-quality solutions to both catastrophic forgetting and ensuring appropriate capacity, and therefore make GPs an excellent model for studying continual learning.
@@
However, developing a practical continual learning in GPs is not as straightforward as it is in finite dimensional linear models, because (for $N$ datapoints) the posterior requires 1) an intractable $O(N^{3})$ computation to compute it exactly, and 2) storing the full training dataset, which breaks the requirements of continual learning. Sparse variational inducing variable methods solve these problems (Titsias, 2009), by introducing a small number of $M$ inducing points that control the capacity of the posterior approximation. In certain settings (e.g. bounded inputs), this approximation is near-exact even when $M\ll N$ (Burt et al., 2019, 2020). This property has allowed continual learning methods to be developed for GPs that perform very closely to full-batch methods (Bui et al., 2017; Maddox et al., 2021; Chang et al., 2023a), provided $M$ is large enough.
@@
As in neural network models, automatically selecting the capacity $M$ is an open problem, with several proposed solutions. Kapoor et al. (2021) acknowledge the need for scaling the capacity with data size, and propose VAR-GP (Variational Autoregressive GP) which adds a fixed number of inducing points for every batch. However, this number may be too small, leading to poor performance, or too large, leading to wasted computation. Galy-Fajou & Opper (2021) propose OIPS (online inducing point selection), which determines $M$ through a threshold on the correlation with other inducing points, and is currently the only adaptive method for GPs, although we show that it still relies on tuned hyperparameters to work.
@@
In this work, we propose to select the capacity of the variational approximation by selecting an appropriate tolerance in the KL gap to the true posterior. This criterion works within the same computational constraints as existing GPs continual learning methods, adapts the capacity to the dataset to minimise computational waste while retaining near-optimal predictive performance. Our method has a single hyperparameter that we keep fixed to a single value, which produces similar trade-offs between size and performance across benchmark datasets with different characteristics.
@@ 3.1 Sparse Variational Gaussian Processes
We consider the typical regression setting, with training data consisting of $N$ input/output pairs $\{\mathbf{x}_{n},y_{n}\}_{n=1}^{N},\mathbf{x}_{n}\in\mathbb{R}^{D},y_{n}\in\mathbb{R}$. We model $y_{n}$ by passing $\mathbf{x}_{n}$ through a function followed by additive Gaussian noise $y_{n}=f(\mathbf{x}_{n})+\epsilon_{n},\epsilon_{n}\sim\mathcal{N}(0,\sigma^{2})$, and take a Gaussian process prior on $f\sim\mathcal{GP}(0,k_{\theta}(\cdot,\cdot))$ with zero mean, and a kernel $k$ with hyperparameters $\theta$. We wish to find the posterior (for prediction) and marginal likelihood (for finding $\theta$). We need approximations, despite these quantities being available in closed form (Rasmussen & Williams, 2005), since they have a computational cost of $O(N^{3})$ that is too high, and require all training data (or statistics greater in size) to be stored, both of which are prohibitive for continual learning. Variational inference requires a lower $O(NM^{2})$ computational and $O(NM)$ memory cost by selecting an approximation from a set of tractable posteriors:
@@
$=\mathcal{N}\left(f(\cdot);\mathbf{k}_{\cdot\mathbf{u}}\mathbf{K}_{\bf uu}^{-1}\mathbf{m},\right.$
$\quad\left.k(\cdot,\cdot)-\mathbf{k}_{\cdot\mathbf{u}}\mathbf{K}_{\bf uu}^{-1}(\mathbf{K}_{\bf uu}-\mathbf{S})\mathbf{K}_{\bf uu}^{-1}\mathbf{k}_{\mathbf{u}\cdot}\right)\,,$ (2)
with $[\mathbf{K}_{\bf uu}]_{ij}=k(\mathbf{z}_{i},\mathbf{z}_{j})$, $[\mathbf{k}_{\cdot\mathbf{u}}]_{i}=[\mathbf{k}_{\mathbf{u}\cdot}^{T}]_{i}=k(\cdot,\mathbf{z}_{i}),\mathbf{Z}=\{\mathbf{z}_{m}\}_{m=1}^{M}$, and $q(\mathbf{u})=\mathcal{N}(\mathbf{u};\mathbf{m},\mathbf{S})$. The variational parameters $\mathbf{m},\mathbf{S},\mathbf{Z}$ and hyperparameters $\theta$ are selected by maximising the Evidence Lower Bound (ELBO) (Hensman et al., 2013). This simultaneously minimises the KL gap $\mathrm{KL}[q(f)\,||\,p(f|\mathbf{y},\theta)]$ between the approximate and true GP posteriors (Matthews et al., 2016; Matthews, 2017), and maximises an approximation to the marginal likelihood of the hyperparameters:
$\mathcal{L}_{\text{ELBO}}=\sum_{i=1}^{N}\mathbb{E}_{q(f(\mathbf{x}_{i}))}[\log p(y_{i}|f(\mathbf{x}_{i}),\theta)]\\ -\mathrm{KL}\left[q(\mathbf{u})\,\|\,p(\mathbf{u}|\theta)\right]\;.$ (3)
The variational approximation has the desirable properties (van der Wilk, 2019) of 1) providing a measure of discrepancy between the finite capacity approximation, and the true infinite capacity model, 2) arbitrarily accurate approximations if enough capacity is added (Burt et al., 2019, 2020), and 3) retaining the uncertainty quantification over the infinite number of basis functions. In this work, we will particularly rely on being able to measure the quality of the approximation to help determine how large $M$ should be.
@@ 3.2 Sparse Gaussian Processes are Equivalent to Single-Layer Neural Networks
The equations of the predictor (eq. (2)) show the strong connection between Sparse Gaussian processes and neural networks. The kernel $k(\cdot,z_{i})$ forms the nonlinearity, with the weights being parameterised as $\mathbf{K}_{\bf uu}^{-1}\mathbf{m}$. For inner product kernels $k(\mathbf{x},\mathbf{Z})=\sigma(\mathbf{Z}\mathbf{x})$ like the arc-cosine kernel (Cho & Saul, 2009), the connection becomes stronger, with $\mathbf{Z}$ as the input weights. This construction also arises from other combinations of kernels and inter-domain inducing variables (Dutordoir et al., 2020; Sun et al., 2020), and has also showed equivalences between deep Gaussian processes and deep neural networks with activations similar to the ReLU (Dutordoir et al., 2021). As a consequence, our method for determining the number of inducing variables in a sparse GP, equivalently finds the number of neurons needed in a single-layer neural network. As such, we hope the ideas presented in this work can inspire similar mechanisms for adaptive neural networks.
@@ 3.3 Online Sparse Gaussian Processes
In this work, we use the extension of the sparse variational GP approximation to the continual learning case developed by Bui et al. (2017). Here, we present a modified derivation that clarifies 1) how the online ELBO provides an estimate to the full-batch ELBO, and 2) when this approximation is accurate.
We update our posterior and hyperparameter approximations after each batch of new data $\{\mathbf{X}_{n},\mathbf{y}_{n}\}$. While we do not have access to data from older batches $\{\mathbf{X}_{o},\mathbf{y}_{o}\}$, the parameters specifying the approximate posterior $q_{o}(f)=p(f_{\neq\mathbf{a}}|\mathbf{a},\theta_{o})q_{o}(\mathbf{a})$ are passed on. We construct this approximate posterior as in eq. (1) but with $\mathbf{a}=f(\mathbf{Z}_{o})$ and the old hyperparameters $\theta_{o}$. Given the “old” $q_{o}(f)$, online sparse GPs construct a “new” approximation $q_{n}(f)$ of the posterior for all observed data $p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})$, which can be written as:
$p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})$ $=\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)p(\mathbf{y}_{o}|f)}{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}$
$=\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)}{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}\frac{p(f|\mathbf{y}_{o},\theta_{o})p(\mathbf{y}_{o}|\theta_{o})}{p(f|\theta_{o})}\;.$
We denote the new variational distribution as $q_{n}(f)=p(f_{\neq\mathbf{b}}|\mathbf{b},\theta_{n})q_{n}(\mathbf{b})$ where $\mathbf{b}=f(\mathbf{Z}_{n})$ and $\theta_{n}$ is the new hyperparameter setting. The KL divergence between the exact and approximate posterior at the current batch is given by:
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$
$=\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}$
$\qquad-\int q_{n}(f)\log\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)p(f|\mathbf{y}_{o},\theta_{o})}{q_{n}(f)p(f|\theta_{o})}\text{d}f\;.$
@@
$\operatorname{KL}[q_{n}(f)\mid\mid p(f|\mathbf{y}_{o},\mathbf{y}_{n},\theta_{n})]$ (4)
$=\log\frac{p(\mathbf{y}_{n},\mathbf{y}_{o}|\theta_{n})}{p(\mathbf{y}_{o}|\theta_{o})}$
$\quad-\underbrace{\int q_{n}(f)\log\frac{p(f|\theta_{n})p(\mathbf{y}_{n}|f)q_{o}(f)}{q_{n}(f)p(f|\theta_{o})}\text{d}f}_{=\hat{\mathcal{L}}}+\Phi$
where $\Phi=-\int q_{n}(f)\log\frac{p(f|\mathbf{y}_{o},\theta_{o})}{q_{o}(f)}\text{d}f$. We cannot compute $\Phi$ due to its dependence on the exact posterior, so we drop it and use the remaining term as our “online ELBO” training objective, which simplifies as:
^$\widehat{\mathcal{L}}:=\int q_{n}(f)\left[\log\frac{p(\mathbf{b}|\theta_{n})q_{o}(\mathbf{a})p(\mathbf{y}_{n}|f)}{q_{n}(\mathbf{b})p(\mathbf{a}|\theta_{o})}\right]\text{d}f,$ (5)
since $q_{o}(f)=p\left(f_{\neq\mathbf{a}}\mid\mathbf{a},\theta_{o}\right)q_{o}(\mathbf{a})$ and similarly for $q_{n}(f)$.
Maximising $\widehat{\mathcal{L}}$ will accurately minimise the KL to the true posterior when $\Phi$ is small, which is the case when the old approximation is accurate, i.e. $q_{o}(f)\approx p(f|\mathbf{y}_{o},\theta_{o})$ for all values of $f(X_{o})$ (with $\Phi=0$ in the case of equality). In our continual learning procedure, we will keep our sequence of approximations accurate by ensuring they all have enough inducing points, which will keep $\Phi$ small.
To get our final bound, we perform a change of variables for the variational distribution $q_{o}(\mathbf{a})=\mathcal{N}(\mathbf{a};\mathbf{m}_{\mathbf{a}},\mathbf{S}_{\mathbf{a}})$ to use the likelihood parametrisation (Panos et al., 2018):
$q_{o}(\mathbf{a})$ $=\frac{\mathcal{N}(\mathbf{a};\tilde{\mathbf{m}}_{\mathbf{a}},\mathbf{D}_{\mathbf{a}})\mathcal{N}(\mathbf{a};0,\mathbf{K}^{\prime}_{\mathbf{aa}})}{\int\mathcal{N}(\mathbf{a};\tilde{\mathbf{m}}_{\mathbf{a}},\mathbf{D}_{\mathbf{a}})\mathcal{N}(\mathbf{a};0,\mathbf{K}^{\prime}_{\mathbf{aa}})\text{d}\mathbf{a}}$ (6)
$=\frac{l(\mathbf{a})p(\mathbf{a}\mid\theta_{o})}{\mathcal{N}(\mathbf{a};0,\mathbf{D}_{\mathbf{a}}+\mathbf{K}^{\prime}_{\mathbf{aa}})},$